14. (7 points) The threshold frequency of a certain metal is 6.53 x 1014 Hz. If light whose wavelength is 405 nm strikes theplease list all steps in solving this problem. Thank YOU
Answer #1

When incident a clean mtallic sunface e Hhene a place between electnon of me tal an collisim taker pho forned A pont nengy iaNow, kE=h-4 whane, C=speid light -4 Me mas 9.1M1031 -4 h=6.6610J -24 me 2 a = 405 nm =405A10-9 m he -4 9.IMo 3 C.GR67103 308If you still have any query feel free to ask in the comments and if you like my work please give a thumbs up.